Title: 2 Blazars
Session: Feb 1997 (session 1)
Frequency: 6 cm
Obs time (UT): EU 21:00(19/02)-09:00(20/02)
EU Mk3 Tapes/St.: 3
To give feedback on this experiment, click here
Effelsberg (dgraham at Thu Feb 20 02:56:00 EST 1997) Minor failures
stop for wind 2130 to 0015
Medicina (Operator \(Mariotti\) at Thu Feb 20 04:48:20 EST 1997) Success
It seem everything OK
Westerbork (Tony Foley at Thu Feb 20 05:27:04 EST 1997) Success
No known problems
Onsala (Antonis Polatidis at Sun Feb 23 08:41:44 EST 1997) Minor failures
Due to high winds (>14m/s) we stowed the antenna between UT 2310 and Ut 0558 and then from 0635 to the end of the schedule. No known problems for the recorded data.
Shanghai (Huang Xinyong at Wed Mar 5 04:52:02 EST 1997) Minor failures
Clock offset = 0.06 msec. Tape vacuum failed four times in first
scan(21:00:00-21:13:00). Weather fine.
Jodrell (Alastair Gunn at Wed Mar 5 07:29:01 EST 1997) Minor failures
Telescope was parked due to high winds until 03:40 UT. No other
known problems. Tapes IAA00326,1CD3 and MPIR1135,D71B sent to Bonn
via Woodcroft on 25/02/97.
Cambridge (Alastair Gunn at Wed Mar 5 07:43:15 EST 1997) Minor failures
Field system crashes occurred at 03:48, 06:30 and 07:52. Telescope
was parked due to high winds until about 04:00 UT. No other known
problems. Tapes IAA00141,40FB and EVN00023,B2AC sent to Bonn via
Woodcroft on 24/02/97.
Calibration (Fredrik T Rantakyro at Thu May 8 11:33:35 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the
Feb97 directory with the name em21a.antab. This file contains the
preliminary calibration results and may be updated when I receive
updated information from the telescopes and the results from the
datareduction of the calibration experiments.
MPIfR (Walter Alef at Wed Jan 7 11:57:26 MET 1998) Minor failures
Crimea and Noto did not observe. No fringes with Urumqi. Several antennas lost
time due to high winds. More than half of telescopes lost on average. Problems
with recording quality.